MsBrains (ms-brains):

Please help me three questions, All you have to do is check my work. I will fan and medal!!

MsBrains (ms-brains):

I think the answer is D

MsBrains (ms-brains):

@DN_20

MsBrains (ms-brains):

Are you going to help @mathstudent55

OpenStudy (dn_20):

So 4(a+3) = 4a +12 replace a = x, which x = any real #s So 4(x+3) = 4x +12 So d is correct.